### v2

# OmniRace: 6D Hand Pose Estimation for Intuitive Guidance of Racing Drone

Abstract
This paper presents the OmniRace approach to controlling a racing drone with 6-degree of freedom (DoF) hand pose estimation and gesture recognition. To our knowledge, it is the first-ever technology that allows for low-level control of high-speed drones using gestures. OmniRace employs a gesture interface based on computer vision and a deep neural network to estimate a 6-DoF hand pose. The advanced machine learning algorithm robustly interprets human gestures, allowing users to control drone motion intuitively. Real-time control of a racing drone demonstrates the effectiveness of the system, validating its potential to revolutionize drone racing and other applications. Experimental results conducted in the Gazebo simulation environment revealed that OmniRace allows the users to complite the UAV race track significantly (by 25.1%) faster and to decrease the length of the test drone path (from 102.9 to 83.7 m). Users preferred the gesture interface for attractiveness (1.57 UEQ score), hedonic quality (1.56 UEQ score), and lower perceived temporal demand (32.0 score in NASA-TLX), while noting the high efficiency (0.75 UEQ score) and low physical demand (19.0 score in NASA-TLX) of the baseline remote controller. The deep neural network attains an average accuracy of 99.75% when applied to both normalized datasets and raw datasets. OmniRace can potentially change the way humans interact with and navigate racing drones in dynamic and complex environments. The source code is available at https://github.com/SerValera/OmniRace.git.
I Introduction
Figure 1: (a) Visual estimation of human hand pose in 6-DoF and gesture recognition. (b) Experimental setup for the evaluation of the gesture control interface. (c) Drone race track in a simulation environment for the evaluation of the control interfaces.
Drone racing from a first-person perspective has been extensively explored due to the challenges it poses, such as adaptive control, navigation in cluttered environment, and human decision-making [1]. In this scenario, pilots navigate the race track via a video transmitted from an onboard camera. The drone’s ability to move with six degrees of freedom within three-dimensional space requires pilots to have advanced visual-motor coordination skills to successfully complete the race within the shortest possible time. The difficulty in controlling drones is caused by a multitude of factors, including the high velocity of the drone’s flight, the requirement for precise control to avoid obstacles, and the cognitive ability to process visual feedback from the drone’s perspective. Learning to skillfully control a racing drone is not a simple task. It demands extensive practice over a prolonged period. However, the duration varies significantly among individuals, with estimates ranging from several weeks to months, depending on the frequency of practice, as well as individual’s inherent motor abilities [2].
The concept of controlling drones through hand motion presents an immersive and intuitive approach in the field of robotics. This method emulates a natural human-machine interface, allowing individuals to navigate robots in a manner akin to the manipulation of physical objects. While this technique may not provide the superlative precision offered by traditional control systems, it introduces a fascinating area of research that explores the integration of human motor skills with robotic control. This form of interaction is reminiscent of the Jedi’s use of “The Force” in the popular Star Wars franchise [3], a notion that has captivated audiences worldwide. The allure lies in the seemingly magical ability to influence one’s surroundings through mere gestures, creating a sense of empowerment and control. This paper aims to explore this fascinating concept and explore the potential of using this “force” to control drones through human hand movements. We explored how this innovative technology works, its potential applications, and the future it holds for drone racing.
In this work we presented a system for controlling a drone using just gestures without any additional wearable devices or controllers. We built on our previous work, SwarmPaint [4], where we introduced the concept of a swarm control interface that allows an operator to guide the swarm through path drawing and formation control. Our current research presents an upgraded version of the drone guiding interface. In addition to high-level commands, the OmniRace interface allows low-level control of the drone’s position by setting the orientation and throttle for the drone’s autopilot. This allows the user to understand the current orientation of the drone and to perform more complex maneuvers with just hand movements. In addition, we tested our gesture interface and compared it against a baseline remote controller on a racing track in a simulation environment.
II Related Work
II-A Human-Robot Interaction
The field of Human-Robot Interaction (HRI) has seen significant advancements in recent years [5], particularly in the area of gesture-based control [6]. Cutting-edge studies have been conducted, resulting in the development of cognitive human-machine interfaces and interactions that support adaptive automation in one-to-many applications, where a single human operator monitors and coordinates multiple Unmanned Aerial Vehicles (UAV) [7]. User-case studies conducted in the field of HRI demonstrate that drones are being actively utilized in real world scenarios. For example, HRI research is aiming at medicinal delivery [8] or forest firefighting missions [9]. This necessitates a working environment where humans and robots coexist and have the capability to effectively control and interact with drones. The human-drone interaction aspect of this research is crucial to ensure these technologies are user-friendly, safe, and efficient in real-world scenarios. For example, HyperPalm [10] is a novel gesture interface for intuitive human-robot interaction with quadruped robots. DroneARchery [11] is an augmented reality human-drone interaction driven by RL-based swarm behavior. Despite these advancements, there are still many challenges to be addressed, such as improving the robustness of gesture recognition in diverse and changing environments [12], reducing the computational requirements of machine learning (ML) and artificial intelligence (AI) algorithms [13], and enhancing the safety and reliability of gesture-based drone control [14].
II-B Sensors and Devices for Gesture Detection
The technology relies on sensors and devices that can detect human hand movements and gestures. These devices include but are not limited to IMU, depth cameras, and infrared sensors. Accelerometers and gyroscopes [15] are often embedded in wearable devices to detect hand motion and to tackle muscle sensors [16], while depth cameras [17] and infrared sensors [18] are used to capture hand shape and gestures in 3D. Also the combination of the sensor can provide advance control of the swarm of drones [19]. These sensors provide raw data which is then processed and interpreted as control commands for the drone. For instance, a specific hand gesture could be interpreted as a command to move the drone forward, backward, left, or right [20]. Wearable devices provide a large amount of data, however, they reduce mobility and ease of control. Visual-based sensors [21] are most actively used in human hand gesture detection for several reasons: they offer high accuracy, are non-intrusive, allow for real-time interaction, and can be integrated with Machine Learning and Artificial Intelligence. Most of the proposed control methods provide the opportunity to control a drone at a high level of commands, which does not allow for quick maneuvering in conditions of drone racing at high speeds.
II-C ML and AI Methods for Gesture Recognition
ML and AI play crucial roles in the interpretation of sensor data into meaningful commands. ML algorithms are used to train models that can recognize specific hand gestures based on sensor data [22]. These algorithms can learn from a set of training data containing various hand gestures and their corresponding drone commands, allowing them to accurately identify these gestures in real-time. On the other hand, AI techniques such as neural networks [23] and deep learning [24] are used to improve the accuracy and efficiency of gesture recognition. These techniques allow the system to learn complex patterns and relationships between different hand gestures and drone commands, enabling it to handle a wider range of gestures and more complex control tasks. Despite these advancements, there are still challenges to be addressed. These include the variability in human gestures due to differences in individual style, speed, and amplitude of gestures, as well as environmental factors such as lighting conditions and background noise [25]. Furthermore, the computational complexity [26] of machine learning and AI algorithms can also pose challenges in real-time gesture recognition applications.
III OmniRace Drone Control Scenario
III-A System Overview
The OmniRace system consists of the following components: computer vision, which includes a hand recognition module using the MediaPipe library, a pre-trained deep neural network (DNN) model for gesture recognition, and an algorithm to estimate the six degrees of freedom (6-DoF) of the operator’s hand position and orientation. The information derived from the hand is then transformed into a control signal for a drone in a software-in-the-loop (SITL) environment, such as a Gazebo simulation or an actual SpeedyBee F405 drone with an ArduPilot autopilot firmware. The SITL gazebo plugin allows for fully simulated drone control, using the MAVROS interface, with the same behavior as a real drone. Input video frames for CV processing are received from a RealSense D435 camera at a resolution of 848 x 480 pixels and a frame rate of 30 fps. A drone remote controller, LiteRadio3 Pro, is used for manual control and has been compared with the proposed gesture control method. All modules are shown in Fig. 2
Figure 2: Algorithm pipeline of the 6-DoF gesture-based race drone control.
III-B DNN-based Gesture Recognition
Figure 3: Structure of the DNN training model. 42 input parameters (coordinates x, y of one-hand landmarks), hidden layer (168, 546), output layer 8 (gestures).
The MediaPipe Hand Landmarker [27] was utilized as an advanced solution for tracking the movements of hands and fingers with high precision. The way it accomplishes this is by deriving 21 specific landmarks (X, Y) from each image frame of a hand. The gesture dataset was gathered from five individuals. This dataset is comprised of 8000 arrays of images, each containing the coordinates of 21 landmarks. This collected dataset additionally includes different hand positions for eight unique gestures. Thus, for every gesture, there is a unique set of 1000 point arrays representing the specific position and motion of the hand. This enables a very detailed and precise representation of each gesture. Overall, OmniRace utilizes eight gestures for high-level drone control (see Fig. 6), which include: “one”, “two”, “three”, “four”, “five”, “okay”, “rock”, and “thumbs up”. The train-test split is 6400 to 1600, representing 80% to 20% data distribution. The DNN model, depicted in Fig. 3, consists of three layers: input, hidden, and output. The input layer receives data features and the output layer delivers predictions of one of the eight gestures after processing through two hidden layers. The accuracy of gesture recognition on the test dataset is remarkably high, standing at 99.75%. Epochs of learning are presented in Fig. 4.
Figure 4: Graphs depicting accuracy (a) and loss (b) functions across multiple epochs. The yellow line represents validation data, while the blue line represents training data of the network learning process.
III-C 6-DoF Hand Pose Estimation and Drone Control
The process of determining the position and orientation of the hand involves calculating the relative translation of a hand in 3D space. The colored images from Fig.  5(a) are used to obtain 2D positions of the 21 landmarks, while the depth frames are utilized to gather data for the 3D positioning of each hand point, as shown in Fig.  5(b). We calculated the hand orientations in 3D by applying the method outlined in [28]. As shown in Fig.  5(a), the hand normal vector $\overrightarrow{Z}$ is calculated as a vector product
$\overrightarrow{Z}=\overrightarrow{AC}\times\overrightarrow{AB}.$ (1)
The hand normal vector $\overrightarrow{X}$ is calculated as a connection vector from the hand point $A$ to mean position $M$ of the hand point $A,B$, and $C$. The equation is shown as
$M=\frac{1}{3}(A+B+C);\qquad\overrightarrow{X}=\overrightarrow{AM}$ (2)
where the vector $\overrightarrow{X}$ is also normalized to a unit vector. The calculation of the hand normal vector $\overrightarrow{Y}$ is ultimately done as the cross product of the normalized vectors $\overrightarrow{X}$ and $\overrightarrow{Z}$. Following this, the orientation of the hand is represented using quaternions $(q_{\omega},q_{x},q_{y},q_{z})$, as illustrated below:
$q_{\omega}=\frac{\sqrt{\overrightarrow{X_{x}}+\overrightarrow{Y_{y}}+\overrightarrow{Z_{z}}}}{2};$ (3)
$q_{x}=\frac{\overrightarrow{Z_{y}}+\overrightarrow{Y_{z}}}{4q_{\omega}};q_{y}=\frac{\overrightarrow{X_{z}}+\overrightarrow{Z_{x}}}{4q_{\omega}};q_{z}=\frac{\overrightarrow{Y_{x}}+\overrightarrow{X_{y}}}{4q_{\omega}}.$ (4)
Then the quaternions are converted to Euler angles and represent as roll, pitch, yaw
$\begin{bmatrix}\phi\\ \theta\\ \psi\end{bmatrix}=\begin{bmatrix}atan2(2(q_{\omega}q_{x}+q_{y}q_{z}),1-2(q_{x}^{2}+q_{y}^{2}))\\ -\pi/2+2atan2(\sqrt{1+2(Q)},\sqrt{1-2(Q})\\ atan2(2(q_{\omega}q_{x}+q_{x}q_{y}),1-2(q_{y}^{2}+q_{z}^{2}))\\ \end{bmatrix}$ (5)
where
$Q=q_{\omega}q_{y}-q_{x}q_{z}.$ (6)
Orientation angles, as shown in Fig.  5(c), namely yaw, pitch, and roll, are transferred from the hand to the drone control system depicted in Fig.  5(d). This allows the operator to set the drone’s orientation at a low control level. The throttle is controlled by the depth value of point $M$.
The eight gestures recognized are utilized as high-level commands for the drone. Using a sequence of gestures allows the operator to send different commands. These commands include actions such as arming or disarming the drone, takeoff, landing, and setting the drone in flight mode or selecting a speed coefficient that is convenient for the user. The “five” is the main gesture to control a drone by orientation. When changing the gesture to select the speed or command, the drone switches to hover mode for safe flight. The set of gesture commands is shown in Fig. 6.
Figure 5: (a) 21 landmarks (x, y) of right-hand joint points detected. (b) Depth image of the hand with identified landmarks. (c) 3D hand pose and orientations calculated using ABCM points from color and depth images. (d) Transformed angles for drone orientation control.
Figure 6: Combination of gesture commands for high-level drone control.
IV Experimental Evaluation
IV-A Research Methodology
In this work, we conducted a peer evaluation of the developed gesture interface through comparison with the baseline remote controller. We invited 10 participants to engage in a drone race in the Gazebo simulation using two interfaces to control the drone flight. The recorded trajectories are shown in Fig. 7. Five participants did not have any prior experience in drone control and were labeled as the Beginner group, while five participants had some prior experience in drone control and were labeled as the Intermediate group. Participants started by controlling the drone initially using a remote controller, then by using the developed OmniRace gesture control interface. We also allowed for short breaks between attempts and when switching control methods. Each participant was presented with a challenge to complete a race consisting of 10 gates, as shown in Fig. 1(c). Prior to the commencement of the race, we provided 30 minutes for the participants to familiarize themselves with both control methods. Subsequently, each participant performed three control flights. The racetrack was designed to test the ability to use all control inputs of a drone (yaw, pitch, roll, throttle), which are determined from gestures or remote control interfaces. Since not all the participants were professional drone race pilots, the drone in the simulation was controlled in stabilize mode during the experiment. After the trials, each participant passed a three-part user survey that consisted of unweighted NASA-TLX (calculated in 100 points from a 21-point scale) [29], SUS [30], and UEQ [31] questionnaires to assess the pragmatic and hedonic qualities of the gesture interface as well as its overall performance.
IV-B Experimental Results
The results of the user study were processed both in one piece and separately for the Beginner and the Intermediate groups. Table I presents the results of evaluation flight trajectories of all categories, and the results of the user questionnaire are represented in Table II.
Figure 7: Set of recorded trajectories by gesture interface (red lines), and by remote controller (blue lines).
TABLE I: Experimental Results of Recorded Flight Trajectories by Gesture Interface and Remote Controller.
Participants Metrics
Gesture
interface
Remote
Controller
Overall Time, s 103.25 129.18
Length flight path, m 83.70 102.90
Average velocity, m/s 0.92 1.13
Beginner level Time, s 135.19 199.43
Length flight path, m 87.88 118.53
Average velocity, m/s 0.67 0.60
Intermediates level Time, s 71.32 58.92
Length flight path, m 79.51 87.26
Average velocity, m/s 1.16 1.65
Upon evaluation, it was observed that the Beginners were able to navigate the track on average one and a half times faster (135.19 s against 199.43 s) using the gesture control interface. In contrast, the Intermediates finished the race a bit faster by using the remote controller. Interestingly, the hand-controlled trajectories resulted in the shortest paths for all participants (mean 83.7 m) compared to those achieved using the remote control (mean 102.9 m). The average speed of the Beginners was the same across both control interfaces. The best time for gesture control was 54.8 s, with the shortest flight path of 71.34 m. The fastest time for remote control was 49.89 s, with the shortest flight path of 70.42 m. Also, the trajectories appeared smoother when controlled via gesture interface control than by a remote controller, as shown in Fig. 7.
[…]
Score
Gesture
interface
Remote
Controller
Wilcoxon,
alpha of 0.05
NASA-TLX
Mental Demand 31.5 56.5 V=7.5, p=0.04
Physical Demand 29.5 19.0 V=10.0, $p$=0.08
Temporal Demand 32.0 49.0 V=6.0, $p$=0.04
Performance 27.7 39.5 V=14.5, $p$=0.62
Effort 32.0 48.0 V=1.5, $p$=0.01
Frustration 16.5 43.5 V=2.5, $p$=0.02
Overall 29.3 42.6 V=4.0, $p$=0.01
UEQ
Attractiveness 1.57 -0.35 V=0.0, $p$=0.002
Perspicuity 1.55 -0.55 V=0.0, $p$=0.007
Efficiency -0.05 0.75 V=11.0, $p$=0.17
Dependability 0.65 0.20 V=19.0, $p$=0.43
Stimulation 1.40 -0.13 V=0.0, $p$=0.012
Novelty 1.73 -1.80 V=1.0, $p$=0.003
Hedonic Quality 1.56 -0.96 V=1.0, $p$=0.003
Pragmatic Quality 0.717 0.131 V=12.0, $p$=0.13
SUS 78.8 53.5 N/A
Given that the data were non-parametric, a Wilcoxon signed-rank test for paired samples. The NASA-TLX test, as represented in Fig. 8 revealed that participants‘ perceived workload differed significantly between the interfaces (V = 4.0, $p$ = $0.01<0.05$), resulting in a higher perceived workload in the case of the remote controller (M = 42.6, SD = 11.8) compared to the gesture interface (M = 29.3, SD = 9.8). Users expectantly evaluated the physical demand higher for the gesture interface (M = 29.5, SD = 19.2), however, statistical significance (V = 10.0, $p$ = 0.08) was not demonstrated for this metric. On the other hand, frustration (V = 2.5, $p$ = 0.02) was perceived by respondents lower for the gesture interface (M = 16.5, SD = 17.9) compared to the remote control (M = 43.5, SD = 26.8). We evaluated the SUS score as a standard metric for calculating the relative usability of the UAV control interfaces. Gesture Interface nearly reached an “excellent” SUS score of M = 78.8 (SD = 7.47, A- score), followed by Remote Controller which yielded an average SUS score of “OK” with M = 53.5 (SD = 17.00, D score). The UEQ dimensions for all drone control interfaces are visualized in Fig. 9.
Figure 8: The visualization shows all dimensions of the NASA-TLX questionnaire. Black error bars denote 95% confidence interval (CI).
Figure 9: The visualization shows all dimensions of the UEQ questionnaire. Black error bars denote 95% CI. Dotted red lines indicate UEQ’s 0.8 threshold.
In the UEQ evaluation, Gesture Interface received a positive user experience evaluation ($>$ 0.8) for attractiveness, perspicuity, stimulation, and novelty. Most of the metrics for Remote Controller including attractiveness, efficiency, stimulation, and dependability, were scored as neutral (in the range of -0.8 and 0.8), while the highest score was received for efficiency. This result corresponds with the overall high performance of the experienced drone pilots with the controller, although the acquired data did not show a statistical difference in this metric. The results showed lower deviations in hedonic quality metrics, while pragmatic quality did not show a significant difference. One possible reason could be that all groups of users were satisfied more with their accomplishments when personally controlling the drone by hand rather than using the remote controller.
V Conclusions and Future Work
In conclusion, this research presents a novel approach to controlling race drones using a DNN-based 6-DoF pose estimation and hand gesture recognition. The experimental results demonstrate that OmniRace enabled users to navigate the UAV racing track significantly faster, with a 25.1% reduction in completion time, and to shorten the length of the drone’s path during testing by an average of 18% compared to using the usual remote controller. Based on the evaluation results, we found out that the gesture interface was perceived as less mentally demanding (31.5 NASA-TLX score), significantly less frustrating (16.5 NASA-TLX score), and superior in both hedonic and pragmatic qualities (1.56 and 0.717 UEQ scores) compared with the baseline remote controller. At the same time, the experienced pilots were able to achieve a shorter trajectory by 9% when operating through OmniRace interface. This demonstrates the potential of the developed technology to transform drone racing and other applications where intuitive, real-time control of drones is required.
In the future research we will conduct experiments with gesture-based control of real drones. We aim to address the limitations related to camera noise, frame rate, and image resolution to enhance control speed. Additionally, OmniRace can potentially be used to control the racing of drone swarms, which cannot be possible with standard remote controllers.
Acknowledgements
Research reported in this publication was financially supported by the RSF grant No. 24-41-02039.